40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16902     Entry time: Wed Jun 8 16:57:46 2022
Author: Koji 
Type: Update 
Category: IOO 
Subject: Power Outage 220608: IMC restored 

The IMC alignment was restored and the IMC is nicely locking.

Once the vacuum level recovered P<1e-4 torr, the PSL shutter was able to be opened.

The IMC was still flashing, so the lock to TEM00 was possible.

Once it was locked, the MC2 alignment was tweaked and the autolocker and the WFS kicked in to help the locking/alignment.

The transmission is ~13k and seems reasonable considering the low PMC transmission of the PMC (0.672)

ELOG V3.1.3-